OpenStudy (anonymous):

find limit of... limt x-->1 [2x](x-1).....and points of continuity of f(x)=x-[x], for all x belongs to real no..and please what does it mean by [x]?

OpenStudy (anonymous):

i bet it is greatest integer function

OpenStudy (anonymous):

limit should be 2

OpenStudy (anonymous):

nevermind...errr

OpenStudy (anonymous):

\[[x]\] is the greatest integer less than or equal to x. so you have \[[\pi]=3\] and \[[-\pi]=-4\]

OpenStudy (anonymous):

here is a nice picture of \[f(x)=x-[x]\] http://www.wolframalpha.com/input/?i=x-+greatest+integer%28x%29

OpenStudy (anonymous):

discontinuous at the integers

OpenStudy (anonymous):

i need a sol problem and i have ashed 2 different question along with the meaning of [x]..please help

OpenStudy (anonymous):

ok here is a picture of \[[x]\] by itself http://www.wolframalpha.com/input/?i=greatest+integer+%28x%29

OpenStudy (anonymous):

it is called a "step function" because the graph looks like steps.

OpenStudy (anonymous):

it means take the biggest integer less than or equal to x. for example, \[[2.3]=2\] \[[7.999]=7\] \[[-2.4]=-3\] in each case the output is the largest integer less than or equal to the input

OpenStudy (anonymous):

it is discontinuous at the integers because it jumps there. it is constant on the interval \[[n, n+1)\] where n is an integer.

OpenStudy (anonymous):

so for example on the interval \[[3,4)\] it is always 3

OpenStudy (anonymous):

let discuss first question...lim x ->1 [2x](x-1)..so what i have to do...i will have to put 1 or some intiger greater than 1

OpenStudy (anonymous):

this is \[[2x](x-1)\] right?

OpenStudy (anonymous):

you have to take the limit from the right and from the left and see if you get the same thing or not.

OpenStudy (anonymous):

yes...i have to calculate just limit for that function,,no need of right or left

OpenStudy (anonymous):

if 1<x then \[[2x]=2\] so you get \[2(x-1)\] which goes to zero as x goes to 1

OpenStudy (anonymous):

oh but that is the whole point, you have to see if they match up!

OpenStudy (amistre64):

limit only exists if its the same number from the right as it is from the left ....

OpenStudy (anonymous):

what amistre said. \[[x]\] is tricky at integers. it has a jump there so your real job is to check limit from right and left and see if they agree.

OpenStudy (anonymous):

now come to 2ns question f(x)= x-[x] for all x belongs to real no... and in that question we have to find points of discontinuity

OpenStudy (anonymous):

they will match because if \[0<x<1\] you have \[[2x]=1\] so your function is \[x-1\] which also has a limit of zero as x goes to 1

OpenStudy (anonymous):

i sent you a picture of the second problem. you can see it is discontinuous at all integers

OpenStudy (anonymous):

i need a solved problem.. i am unable to understand from pic

OpenStudy (anonymous):

what does "discontinuous" mean? it means there is a jump in the graph, so picture gives it. but lets use some algebra if that would help

OpenStudy (anonymous):

yes solve it by algebra

OpenStudy (anonymous):

lets look at \[x-[x]\] on the interval [0,2) for \[0\leq x <1\] you have \[[x]=0\]

OpenStudy (anonymous):

so on the interval \[[0,1)\] you have \[f(x)=x-[x]=x-0=x\]

OpenStudy (anonymous):

on the interval \[[1,2)\] you have \[[x]=1\] so on that interval \[f(x)=x-[x]=x-1\]

OpenStudy (anonymous):

and so one. so on the interval \[[2,3)\] you have \[f(x)=x-2\] etc etc

OpenStudy (anonymous):

in other words you have a piece wise function. on the interval \[[n,n+1)\] you have \[f(x)=x-n\]

OpenStudy (anonymous):

this is clearly not continuous for any integer n because the limit from the left and from the right do not agree

OpenStudy (anonymous):

ok.thanks

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

if x < n you have \[\lim_{x\rightarrow n^-}f(x)=\lim_{x\rightarrow n-} x-(n-1)=1\] whereas if

OpenStudy (anonymous):

now what will be the limit of x->0 x[1/x]

OpenStudy (anonymous):

and if x > n you have \[\lim_{x\rightarrow n^+}f(x)=\lim_{x\rightarrow n+} x-n=0\]

OpenStudy (anonymous):

there all gory details worked out

OpenStudy (anonymous):

you want a guess? i would say \[\lim_{x\rightarrow 0}x[\frac{1}{x}]=1\]

OpenStudy (anonymous):

yes...but i think it should be zero as we have to put 1 in [1/x] and the it will be x and the put 0?

OpenStudy (anonymous):

but guess what you have to do? take the limit from the left and from the right. that is the whole job here i am sure. see that you get one in any case

OpenStudy (anonymous):

you lost me. you are taking the limit as x goes to zero, not as x goes to one

OpenStudy (anonymous):

so limit goes to zero ... as x=0 xpression will be 0[1/0]..

OpenStudy (anonymous):

this is what you have to check, that \[\lim_{x\rightarrow 0^+}x[\frac{1}{x}]=\lim_{x\rightarrow 0^-}x[\frac{1}{x}]\] and my guess is that they do, and that you get 1 in both cases

OpenStudy (anonymous):

by putting 0 both goes to infinity

OpenStudy (amistre64):

another good point to bring out is that the value of the limit does not care about what the actual value of the function is at the point.

OpenStudy (anonymous):

you cannot compute this limit by replacing x by 0 because you will get 0/0 which is meaningless. you have to do something else. also you have to get rid of the greatest integer symbol and write the piecewise version of this thing

OpenStudy (anonymous):

right, what amistre said. this thing has no value if you replace x by zero. you have to find the limit.

OpenStudy (anonymous):

ok ..thank u all..take care ...:)

OpenStudy (anonymous):

good luck. btw i am sure the answer is 1